# Bayesian Modular Inference for Copula Models with Potentially Misspecified Marginals

### 12 Mar 2026 v1

@@ 3.2 Multiple cuts of marginals
Integrating over $u$ gives a pseudo likelihood where the marginals for $D(\delta)$ only use their rank data; see Appendix A. This resulting pseudo likelihood is a mixed density, so that we call it a mixed pseudo likelihood. When $\delta=(0,\ldots,0)^{\top}$, $D^{c}(\delta)=\emptyset$ and (3.2) is exactly equal to (8), whereas when $\delta=(1,\ldots,1)^{\top}$, $D(\delta)=\emptyset$ and (3.2) is the copula density evaluated at points $u_{ij}=F_{j}(y_{ij};\eta_{j})$ for all $i,j$.
@@ 4 Theoretical behavior of SMI copulas
4.1 Influence parameter impact
The behavior of the SMI posterior at (14) depends on the value of the influence parameter $\gamma$. This is easy to characterize in the limiting cases where $\gamma_{j}\in\{0,1\}$ for each $j=1,\dots,d$. For example, when $\gamma=(1,\ldots,1)^{\top}$, as outlined above $p_{\text{SMI},1}(\eta,\psi\mid\mathcal{D})=p(\eta|\psi,\mathcal{D})$ is the conventional posterior. Whereas, when $\gamma=(0,\ldots,0)^{\top}$,
$p_{\text{SMI},\gamma}(\psi,u,\eta,\widetilde{\eta}\mid\mathcal{D})=p_{\text{SMI},\gamma}(\psi,u,\eta\mid\mathcal{D})\propto\underbrace{p_{\mathrm{pl}}(r(\mathcal{D}),u\mid\psi)p(\psi)}_{p_{\mathrm{cut}}(\varphi,u\mid r(\mathcal{D}))}p(\eta\mid\psi,\mathcal{D}),$
so that influence from $\widetilde{\eta}$ onto $\psi$ is fully cut, and the SMI posterior for is the cut posterior at (7).
More generally, let $D(\gamma)=\{j;\gamma_{j}=0\}$ denote the indicies of the fully cut marginals, and $D^{c}(\gamma)=\{j;0<\gamma_{j}\leq 1\}$ denotes the indices of non-zero influence parameters, with $\text{Card}[D^{c}(\gamma)]>1$, then the SMI posterior is a non-linear combination of the cut posterior (7) for the indices where $j\in D(\gamma)$, and a semi-modular posterior for the remaining components. As in Section 3.2, we partition $\widetilde{\eta}$ into full-cut $\widetilde{\eta}_{D(\gamma)}$ and remaining components $\widetilde{\eta}_{D^{c}(\gamma)}$. Then,
$\int p_{\text{SMI},\gamma}(\psi,u,\eta,\widetilde{\eta}\mid\mathcal{D})\mathrm{d}\widetilde{\eta}$
$\propto\int p(\psi)p_{\text{SMI},\gamma}(\mathcal{D},u\mid\psi,\widetilde{\eta}_{D^{c}(\gamma)},\widetilde{\eta}_{D(\gamma)})\prod_{i=1}^{n}\prod_{j\in D^{c}(\gamma)}p(\widetilde{\eta}_{j})^{1/n}f_{j}(y_{ij}\mid\widetilde{\eta}_{j})\prod_{j\in D(\gamma)}p(\widetilde{\eta}_{j})^{1/n}f_{j}(y_{ij}\mid\widetilde{\eta}_{j})\mathrm{d}\widetilde{\eta}$
$\not\propto\int p_{\text{SMI},\gamma}(\mathcal{D},u\mid\psi,\widetilde{\eta}_{D^{c}(\gamma)},0)\prod_{i=1}^{n}\prod_{j\in D^{c}(\gamma)}p(\widetilde{\eta}_{j})^{1/n}f_{j}(y_{ij}\mid\widetilde{\eta}_{j})\mathrm{d}\widetilde{\eta}_{D^{c}(\gamma)}.$
That is, so long as $\gamma_{j}\neq 0$ for even a single index, the influence of the fictional parameters $\widetilde{\eta}$ does not wash out of the semi-modular posterior (SMP): even when $\gamma_{j}=0$ the $j$-th marginal parameter $\widetilde{\eta}_{j}$ still delivers a joint posterior that depends on the fictitious $\widetilde{\eta}_{D(\gamma)}$ through the term
$\int\prod_{j\in D(\gamma)}p(\widetilde{\eta}_{j})^{1/n}f_{j}(y_{ij}\mid\widetilde{\eta}_{j})\mathrm{d}\widetilde{\eta}_{D(\gamma)}.$
@@ 4.2 $\gamma$-dependent concentration results
Our SMPs unique dependence on $\gamma$ ultimately implies that the point onto which the posterior concentrates must be thought of as being $\gamma$-dependent. This result suggests that, unlike the usual learning rate encountered in generalized Bayesian inference, the value of $\gamma$ directly influences not just posterior width but location (i.e., concentration) as well. This fact alone suggests that one should optimize over $\gamma$ to determine which values deliver accurate inferences and predictions.
Given that the SMI posterior contains the fully cut posterior at (7) as a special case, $\gamma_{j}=0$ for all $j$, and given that the theoretical behavior of standard posteriors, $\gamma_{j}=1$ for all $j$, are well-known, we restrict our theoretical analysis to values where $\gamma_{j}\in[0,1)$ and with $\gamma_{j}\neq 0$ for at least some $j$.
@@
$M^{\gamma}_{n}(\varphi):=-\log\int p_{\mathrm{SMI},\gamma}(\mathcal{D}\mid\psi,\widetilde{\eta})\mathrm{d}u,\quad\mathcal{M}(\varphi;\gamma):=\lim_{n\rightarrow\infty}\mathbb{E}[M^{\gamma}_{n}(\varphi)/n].$
Further, we can define the $\gamma$-dependent pseudo-true value onto which the posterior concentrates as
$\varphi_{\star}(\gamma):=\operatornamewithlimits{argmin}_{\varphi\in\Psi\times\mathcal{E}}\mathcal{M}(\varphi;\gamma).$
Since $p_{\mathrm{SMI},\gamma}(\varphi\mid\mathcal{D})$ can be represented as a generalized posterior based on the loss $M_{n}^{\gamma}(\varphi)$, whose corresponding population optimizer $\varphi_{\star}(\gamma)$ changes with $\gamma$, it is clear that the theoretical behavior of $p_{\mathrm{SMI},\gamma}(\varphi\mid\mathcal{D})$ will depend on the value of $\gamma$. The following result clarifies this dependence, and shows that the SMI posterior for $\psi$ concentrates at the standard rate to a $\gamma$-dependent optimization.
@@
Under Assumptions 1-4 in Appendix B, for a positive sequence $r_{n}\rightarrow 0$, $M_{n}$ large enough, and some $C>0$,
$\mathbb{E}\int\mathds{1}\{\|\psi-\psi_{\star}(\gamma)\|>M_{n}r_{n}\}p_{\mathrm{SMI},\gamma}(\psi\mid\mathcal{D})\mathrm{d}\psi\leq C/M_{n}.$
To derive a result for the behavior of $p(\eta\mid\varphi,\mathcal{D})$ we require a few additional definitions: let $\ell_{n}(\varphi)=\prod_{i}c(u_{i1},\dots,u_{id};\varphi)\prod_{j}f(y_{ij}\mid\eta)$, $\mathcal{L}_{\infty}(\varphi)=\lim_{n}\mathbb{E}[\ell_{n}(\varphi)]/n$ and let $\eta_{\star}(\gamma):=\operatornamewithlimits{argmin}_{\eta\in\mathcal{E}}\mathcal{L}_{\infty}\{\eta,\psi_{\star}(\gamma)\}$.
Theorem 2.
Under Assumptions 1-4 in Appendix B, for $M_{n}$ large enough, some $C>0$,
$\mathbb{E}\int\int\mathds{1}\{\|\eta-\eta_{\star}(\gamma)\|>M_{n}r_{n}\}p_{\mathrm{SMI},\gamma}(\eta\mid\psi,\mathcal{D})\mathrm{d}\psi\mathrm{d}\eta\leq C/M_{n}.$
The nature of the SMP construction means that results beyond posterior concentration derived in Theorems 1-2 are unlikely to be enlightening due to the SMPs nonlinear dependence on $\gamma$. In particular, it is not obvious to us that the existing theoretical results on cut posteriors, Frazier and Nott (2025a), are actually satisfied by the SMP, and it appears that additional regularity would be required to obtain results about the calibration of these SMPs. Due to the technical nature of such constructions, we propose to study these issues in a more general setting separately.
@@ 4.3 Interpretation of results
The proposed SMP behaves in a distinctly different fashion to power-posterior SMP approaches: since we have multiple values of $\gamma$, so long as at least one $\gamma_{j}\in(0,1)$, the likelihood for the $j$-th marginal parameter $\widetilde{\eta}_{j}$ is still informed by data, as can be seen by (4.1). Consequently, our theoretical results should be interpreted as showing that the SMP defines a posterior process $\{\gamma\in[\varepsilon,1-\varepsilon]^{d}:p_{\mathrm{SMI},\gamma}(\varphi,\eta\mid\mathcal{D})\}$, for small $\varepsilon>0$, where the points onto which $p_{\mathrm{SMI},\gamma}(\varphi,\eta\mid\mathcal{D})$ concentrate explicitly depend on the values of $\gamma$.
@@
SMPs dependence on $\gamma$ suggests that it may be possible to learn pairs of values $\gamma,\varphi$ according to some measure of “accuracy”. In particular, since each individual value of $\gamma\in[0,1]^{d}$ indexes a generalized posterior, and thus some loss function, we could in-principle select $\gamma$ via the loss function selection method proposed in Jewson and Rossell (2022) where we place uniform priors (on 0 to 1) over each component $\gamma_{j}$. However, in contrast to the analysis in Jewson and Rossell (2022), there is no reason to suspect that the value of $\gamma,\varphi$ should be uniquely identified from the data.
@@
which will clearly depend on $\gamma_{j}$, and will therefore influence the values of $\varphi$ onto which the posterior concentrates.
That being said, this does not mean that one cannot attempt to target some value in the equivalence class of optima. Herein, we have attempted to select such a value of $\gamma$, and ultimately the value of $\varphi_{\star}(\gamma)$, through BO. We note here that the theoretical analysis of such a procedure is complicated by the lack of identification and so we leave a detailed theoretical analysis of the optimal value of $\gamma$, and thus the point onto which the SMP concentrates to further research.
@@ 5 Simulations
The integrals are numerically approximated using Monte Carlo samples from $p_{\text{SMI},\gamma}(\theta\mid\mathcal{D})$ as explained in Appendix C.
@@ Results
Figure 1: Simulations. Average values for $\mathcal{L}_{1}$ (A), $\mathcal{L}_{2}$ (B), $\mathcal{L}_{\text{cop}}$ (C), and $u(\gamma)$ (D) across $m=100$ repetitions for different combinations of $\gamma=(\gamma_{1},\gamma_{2})^{\top}$. Smaller values indicated by lighter colour are preferred. The conventional posterior $\gamma=(1,1)^{\top}$ corresponds to the upper right corner, while the lower left corner is the fully cut posterior $\gamma=(0,0)^{\top}$.
@@ 6.2 Copula model
We assume that the three series $(Y_{\text{VIX},t},Y_{\text{AAA},t},Y_{\text{BBB},t})$ follow a 3-dimensional stationary stochastic process. Because $Y_{t}$ contains lagged values of these three series, this assumption further restricts the parameter space, with $\Omega$ a block symmetric matrix, $\alpha$ constrained, and the marginals for $Y_{j,t}$ and $Y_{j,t-1}$ being the same for each $j\in\{\text{VIX},\text{AAA},\text{BBB}\}$. Full specification of the copula parameterization under this stationary assumption is given in Appendix D, where the dimensions of $\psi$, $\eta$ and $\gamma$ are 15, 12 and 3, respectively. Finally, when evaluating the copula density and function, calculating $F_{X_{j}}^{-1}$ is a key computation, and we do this using the spline interpolation approach outlined in Smith and Maneesoonthorn (2018).
@@ 6.3 Empirical results
which we optimize using BO. This utility function is similar to that discussed in Section 5. Figure 2 shows how the utility (17) improves with the number of SMI posteriors checked. In this example, BO converges fast after a brief exploration of the $\gamma$ space. The fully cut model proposed by Smith et al. (2025) corresponds to $\gamma=(0,0,0)^{\top}$ and has a higher utility value than the conventional posterior where $\gamma=(1,1,1)^{\top}$. The optimal vector of influence parameters maximizing (17) is $\gamma^{\ast}=(1.00,0.61,0.00)^{\top}$.
Figure 2: Convergence of BO over $100$ iterations for the bond yields example. The vertical axis is the utility $u(\gamma)$, and the black line indicates the optimal value $u(\gamma^{\ast})$ against number of $\gamma$ values explored. The values $u((1,1,1)^{\top})$, $u((0,0,0)^{\top})$ are plotted as gray horizontal lines.
@@
Further empirical results, including copula parameter estimates and Kendall’s tau values, are given for all three posteriors in Supporting Information II.
@@ 7 Conclusion
The Appendix contains A) further discussion on the mixed pseudo likelihood (3.2), B) the assumptions for the theoretical analysis presented in Section 4, C) additional information for the simulation study presented in Section 5, and D) the parameterization of the skew-normal copula for the financial application presented in Section 6.
An online Supporting Information contains proofs for the Theorems stated in Section 4, and further results for the financial application not presented in the main text.
Code is publicly available at github.com/kocklucx/CopulaSMI.
Appendix
@@ Appendix A Mixed Pseudo Likelihood
Equation (3.2) in Section 3.2 gives a pseudo likelihood for the multiple cut case that is extended with the auxilary variables $u$. To integrate over these auxilary variables, partition $u_{i}$ into uncut $u_{i,D^{c}(\delta)}=\{u_{ij};j\in D^{c}(\delta)\}$ and cut $u_{i,D(\delta)}=\{u_{ij};j\in D(\delta)\}$ components, and decompose the copula density into $c(u_{i};\psi)=c_{D^{c}}(u_{i,D^{c}(\delta)};\psi)c_{D|D^{c}}(u_{i,D(\delta)}|u_{i,D^{c}(\delta)};\psi)$ where $c_{D^{c}}$ is the marginal copula density of $u_{i,D^{c}(\delta)}$, and $c_{D|D^{c}}$ is the conditional density of $u_{i,D(\delta)}$ given $u_{i,D^{c}(\delta)}$. Then if the indices $D(\delta)=\{j_{1},j_{2},\ldots,j_{r}\}$,
@@
Note that the cut posterior fits the two module case outlined in Section 2.3 by setting $p_{1}(\mathcal{D}|\psi,\eta_{D^{c}(\delta)})=p_{\text{mpl},\delta}(\mathcal{D}|\psi,\eta_{D^{c}(\delta)})\prod_{j\in D^{c}(\delta)}f_{j}(y_{ij};\eta_{j})$ and $p_{2}(\mathcal{D}|\psi,\eta)=p(\mathcal{D}|\psi,\eta)/p_{1}(\mathcal{D}|\psi,\eta_{D^{c}(\delta)})$.
Appendix B Assumptions
[…]
$\mathsf{L}^{\gamma}_{n}(\varphi)=-\frac{1}{n}\left[\log\int p_{\text{SMI},\gamma}(\mathcal{D},u\mid\psi,\widetilde{\eta})\mathrm{d}u+\sum_{i=1}^{n}\sum_{j=1}^{d}\log f_{j}(y_{ij}\mid\widetilde{\eta}_{j})\right].$
Assumption 1.
For some $\varepsilon>0$, $\gamma_{j}\in[\varepsilon,1-\varepsilon]$ for all $j=1,\dots,d$.
Assumption 2.
(i) For any $\varphi\in\Phi$, the function $\varphi\mapsto\mathsf{L}(\varphi;\gamma):=\mathbb{E}\mathsf{L}_{n}^{\gamma}(\varphi)$ exists. (ii) For some $\varepsilon>0$, and any $\gamma\in[0,1-\varepsilon]^{d}$, there exist $\gamma\mapsto\varphi^{\gamma}_{0}$ such that, for any $\delta>0$,
$\inf_{\varphi:\|\varphi-\varphi_{0}(\gamma)\|>\delta}\mathsf{L}(\varphi;\gamma)\geq\mathsf{L}(\varphi^{\gamma}_{0};\gamma).$
(iii) For some $C>0$ and $\alpha>0$, $\|\varphi-\varphi_{0}^{\gamma}\|\leq C[\mathsf{L}(\varphi;\gamma)-\mathsf{L}(\varphi^{\gamma}_{0};\gamma)]^{\alpha}$.
Assumption 3.
There exists a positive sequence $r_{n}\downarrow 0$ as $n\rightarrow\infty$ such that, for any $\gamma\in[\varepsilon,1-\varepsilon]^{d}$,
$\mathbb{E}\left[\sup_{\varphi\in\Phi}|\mathsf{L}_{n}^{\gamma}(\varphi)-\mathbb{E}[\mathsf{L}_{n}^{\gamma}(\varphi)]|\right]\leq r_{n}$
Assumption 4.
Let $\epsilon>0$ with $n\epsilon\geq 1$, and let $\alpha>0$ be as in Assumption 2. For $\mathcal{A}_{\epsilon}:=\{\theta\in\Theta:|\mathsf{L}(\varphi;\gamma)-\mathsf{L}(\varphi^{\gamma}_{0};\gamma)|\leq\epsilon^{\alpha}\}$, the prior satisfies $\int_{\Phi}\mathds{1}\{\varphi\in\mathcal{A}_{\epsilon}\}p(\varphi)\mathrm{d}\varphi\geq e^{-n\epsilon^{\alpha}}$.
A high-level discussion on these assumptions is given in Supporting Information A.
Appendix C Additional information for the simulation study
We consider the three evaluation metrics
Appendix D Parameterization of the skew-normal copula
In Section 6, assume $\widetilde{Y}_{t}=(Y_{\text{VIX},t},Y_{\text{AAA},t},Y_{\text{BBB},t})^{\top}$ follows a stationary stochastic process. Then from Theorem 1 in Smith (2015), the three marginals of $\widetilde{Y}_{t}$ are time invariant, and the copula of $Y_{t}=(\widetilde{Y}_{t}^{\top},\widetilde{Y}_{t-1}^{\top})^{\top}$ is also stationary, in that $C(u_{t},u_{t-1})$ is invariant to $t$. Invariance in the implicit SN copula $C$ follows from invariance in the underlying SN distribution, which can be enforced by the two parameter constraints described below.
$\Omega=\begin{pmatrix}\Omega_{0}&\Omega_{1}\\ \Omega_{1}^{\top}&\Omega_{0}\end{pmatrix}\,$. (18)
The second constraint follows from writing $\alpha=\left(1-\delta^{\top}\Omega^{-1}\delta\right)^{-\frac{1}{2}}\Omega^{-1}\delta$ in terms of the slant parameters $\delta$ in Azzalini and Valle (1996). If we partition $\delta=(\delta_{2}^{\top},\delta_{1}^{\top})^{\top}$ and $\alpha=(\alpha_{2}^{\top},\alpha_{1}^{\top})^{\top}$ consistent with (18), then invariance follows when $\delta_{2}=\delta_{1}$. This implies that $\alpha_{1}=(\Omega_{0}-\Omega_{1})^{-1}(\Omega_{0}-\Omega_{1}^{\top})\alpha_{2}$. All together, $\Omega_{0}$ is parameterized through 3 unique transformed partials, $\Omega_{1}$ is parameterized through 9 unique transformed partials, and $\alpha$ is parameterized by 3 unique values in $\alpha_{2}$, giving an unrestricted copula parameter vector $\psi$ of dimension 15. Because there are only three unique marginals, $\gamma$ is of dimension three.
Supporting Information
Appendix I Theory
This Section contains a discussion on the assumptions, and proofs for the Theorems stated in Section 4 of the manuscript.
I.1 Proofs
Define
$p_{\text{SMI},\gamma}(\psi\mid\mathcal{D})=\int\int p_{\text{SMI},\gamma}(\psi,u,\eta,\widetilde{\eta}\mid\mathcal{D})dud\widetilde{\eta}$

### 02 Apr 2026 v2

@@ 3.2 Multiple cuts of marginals
Integrating over $u$ gives a pseudo likelihood where the marginals for $D(\delta)$ only use their rank data; see Supporting Information A. This resulting pseudo likelihood is a mixed density, so that we call it a mixed pseudo likelihood. When $\delta=(0,\ldots,0)^{\top}$, $D^{c}(\delta)=\emptyset$ and (3.2) is exactly equal to (8), whereas when $\delta=(1,\ldots,1)^{\top}$, $D(\delta)=\emptyset$ and (3.2) is the copula density evaluated at points $u_{ij}=F_{j}(y_{ij};\eta_{j})$ for all $i,j$.
@@ 4.1 $\gamma$-dependent concentration results
Our SMPs dependence on $\gamma$ ultimately implies that the point onto which the posterior concentrates must be viewed as being $\gamma$-dependent. This result suggests that, unlike the usual learning rate encountered in generalized Bayesian inference, the value of $\gamma$ directly influences not just posterior width but location (i.e., concentration) as well. This suggests that one should optimize over $\gamma$ to determine which values deliver accurate inferences and predictions.
Given that the SMI posterior contains the fully cut posterior at (7) as a special case, $\gamma_{j}=0$ for all $j$, and given that the theoretical behavior of standard posteriors, $\gamma_{j}=1$ for all $j$, are well-known, we restrict our theoretical analysis to values where $\gamma_{j}\in[0,1]$ and where $\gamma_{j}\not\in\{0,1\}$ for at least some $j$.
@@
$M_{n}(\varphi;\gamma):=-\log\int p_{\mathrm{SMI},\gamma}(\mathcal{D},u\mid\psi,\widetilde{\eta})\mathrm{d}u,\quad\mathcal{M}(\varphi;\gamma):=\lim_{n\rightarrow\infty}\mathbb{E}[M_{n}(\varphi;\gamma)/n].$
Further, define the $\gamma$-dependent pseudo-true value as
$\varphi_{0}(\gamma):=\operatornamewithlimits{argmin}_{\varphi\in\Psi\times\mathcal{E}}\mathcal{M}(\varphi;\gamma).$
Likewise, for $\ell_{n}(\eta,\psi)=\prod_{i}c(u_{i1},\dots,u_{id};\psi)\prod_{j}f(y_{ij}\mid\eta)$, and $\mathcal{L}_{\infty}(\eta,\psi)=\lim_{n}\mathbb{E}[\ell_{n}(\eta,\psi)]/n$ define
$\eta_{0}(\gamma):=\operatornamewithlimits{argmin}_{\eta\in\mathcal{E}}\mathcal{L}_{\infty}\{\eta,\psi_{0}(\gamma)\}.$
The following result clarifies the SMI posteriors dependence on $\gamma$: posterior concentration occurs at the standard rate but towards a $\gamma$-dependent quantity.
@@
Under Assumptions 1-4 in Supporting Information B, for a positive sequence $r_{n}\rightarrow 0$, $M_{n}$ large enough, some $C>0$ and $\alpha>0$, with probability converging to one,
$\int\mathds{1}\{\|\psi-\psi_{0}(\gamma)\|>CM_{n}^{\alpha}r_{n}^{\alpha}\}p_{\mathrm{SMI},\gamma}(\psi\mid\mathcal{D})\mathrm{d}\psi\rightarrow 0$
and
$\int\int\mathds{1}\{\|\eta-\eta_{0}(\gamma)\|>CM_{n}^{\alpha}r_{n}^{\alpha}\}p_{\mathrm{SMI},\gamma}(\eta,\psi\mid\mathcal{D})\mathrm{d}\psi\mathrm{d}\eta\rightarrow 0.$
The nature of the SMP construction means that results beyond posterior concentration derived in Theorem 1 are unlikely to be enlightening due to the SMPs nonlinear dependence on $\gamma$. In particular, it is not obvious to us that the existing theoretical results on the asymptotic shape of cut posteriors, Frazier and Nott (2025a), would be satisfied for the SMP, and additional regularity conditions would be required to obtain results on the uncertainty quantification for this class of SMPs. Due to the technical nature of such constructions, we propose to study these issues in a more general setting separately.
@@ 4.2 Interpretation of results
The proposed SMP behaves in a distinctly different fashion to power-posterior SMP approaches: since we have multiple values of $\gamma$, so long as at least one $\gamma_{j}\in(0,1)$, the likelihood for the $j$-th marginal parameter $\widetilde{\eta}_{j}$ is still informed by data. Consequently, our theoretical results should be interpreted as showing that the SMP defines a posterior process $\{\gamma\in[0,1]^{d}:p_{\mathrm{SMI},\gamma}(\psi,\eta\mid\mathcal{D})\}$, where the points onto which $p_{\mathrm{SMI},\gamma}(\psi,\eta\mid\mathcal{D})$ concentrate explicitly depend on the values of $\gamma$.
@@
SMPs dependence on $\gamma$ suggests that it may be possible to learn an optimal value of $\gamma$ according to some measure of “accuracy”. In particular, since each individual value of $\gamma\in[0,1]^{d}$ indexes a generalized posterior, and thus some loss function, we could in-principle select $\gamma$ via the loss function selection method proposed in Jewson and Rossell (2022) where we place uniform priors (on 0 to 1) over each component $\gamma_{j}$. However, in contrast to the analysis in Jewson and Rossell (2022), there is no reason to suspect that the value of $\gamma$, and hence $\varphi_{0}(\gamma)$, should be uniquely identified from the data.
@@
which will clearly depend on $\gamma_{j}$, and will therefore influence the values of $\psi,\eta$ onto which the posterior concentrates.
This does not mean that one cannot attempt to target some value in the equivalence class of optima. Herein, we have attempted to select such a value of $\gamma$, and ultimately the value of $\varphi_{\star}(\gamma)$, through BO. We note here that the theoretical analysis of such a procedure is complicated by the lack of identification and so we leave a detailed theoretical analysis of the optimal value of $\gamma$, and thus the point onto which the SMP concentrates to further research.
@@ 5 Simulations
The integrals are numerically approximated using Monte Carlo samples from $p_{\text{SMI},\gamma}(\theta\mid\mathcal{D})$.
@@ Results
Figure 1: Simulations. Average values for $\mathcal{L}_{1}$ (Panel A), $\mathcal{L}_{2}$ (Panel B), $\mathcal{L}_{\text{cop}}$ (Panel C), and $u(\gamma)$ (Panel D) across $m=100$ repetitions for different combinations of $\gamma=(\gamma_{1},\gamma_{2})^{\top}$. Smaller values indicated by lighter colour are preferred. The conventional posterior $\gamma=(1,1)^{\top}$ corresponds to the upper right corner, while the lower left corner is the fully cut posterior $\gamma=(0,0)^{\top}$.
@@ 6.2 Copula model
We assume that the three series $(Y_{\text{VIX},t},Y_{\text{AAA},t},Y_{\text{BBB},t})$ follow a 3-dimensional stationary stochastic process. Because $Y_{t}$ contains lagged values of these three series, this assumption further restricts the parameter space, with $\Omega$ a block symmetric matrix, $\alpha$ constrained, and the marginals for $Y_{j,t}$ and $Y_{j,t-1}$ being the same for each $j\in\{\text{VIX},\text{AAA},\text{BBB}\}$. Full specification of the copula parameterization under this stationary assumption is given in Supporting Information D.1, where the dimensions of $\psi$, $\eta$ and $\gamma$ are 15, 12 and 3, respectively. Finally, when evaluating the copula density and function, calculating $F_{X_{j}}^{-1}$ is a key computation, and we do this using the spline interpolation approach outlined in Smith and Maneesoonthorn (2018).
@@ 6.3 Empirical results
which we optimize using BO. This utility function is similar to that discussed in Section 5. The fully cut model proposed by Smith et al. (2025) corresponds to $\gamma=(0,0,0)^{\top}$ and has a higher utility value than the conventional posterior where $\gamma=(1,1,1)^{\top}$. The optimal vector of influence parameters maximizing (17) is $\gamma^{\ast}=(1.00,0.61,0.00)^{\top}$.
@@
Further empirical results, including copula parameter estimates and Kendall’s tau values, are given for all three posteriors in Supporting Information D.2.
@@ 7 Conclusion
Supporting Information An online Supporting Information contains A) further discussion on the mixed pseudo likelihood (3.2), B) the assumptions and proofs for the theoretical analysis, C) additional information for the simulation study, and D) further results for the financial application (.pdf).
Python code Code is publicly available at github.com/kocklucx/CopulaSMI.
Supporting Information
@@ Appendix A Mixed Pseudo Likelihood
Equation (9) in Section 3.2 gives a pseudo likelihood for the multiple cut case that is extended with the auxilary variables $u$. To integrate over these auxilary variables, partition $u_{i}$ into uncut $u_{i,D^{c}(\delta)}=\{u_{ij};j\in D^{c}(\delta)\}$ and cut $u_{i,D(\delta)}=\{u_{ij};j\in D(\delta)\}$ components, and decompose the copula density into $c(u_{i};\psi)=c_{D^{c}}(u_{i,D^{c}(\delta)};\psi)c_{D|D^{c}}(u_{i,D(\delta)}|u_{i,D^{c}(\delta)};\psi)$ where $c_{D^{c}}$ is the marginal copula density of $u_{i,D^{c}(\delta)}$, and $c_{D|D^{c}}$ is the conditional density of $u_{i,D(\delta)}$ given $u_{i,D^{c}(\delta)}$. Then if the indices $D(\delta)=\{j_{1},j_{2},\ldots,j_{r}\}$,
@@
Note that the cut posterior fits the two module case by setting $p_{1}(\mathcal{D}|\psi,\eta_{D^{c}(\delta)})=p_{\text{mpl},\delta}(\mathcal{D}|\psi,\eta_{D^{c}(\delta)})\prod_{j\in D^{c}(\delta)}f_{j}(y_{ij};\eta_{j})$ and $p_{2}(\mathcal{D}|\psi,\eta)=p(\mathcal{D}|\psi,\eta)/p_{1}(\mathcal{D}|\psi,\eta_{D^{c}(\delta)})$.
Appendix B Theory
This Section contains a discussion on the assumptions, and proofs for the results stated in Section 4 of the manuscript.
B.1 Proofs
Recall
$p_{\text{SMI},\gamma}(\psi\mid\mathcal{D})=\int\int\int p_{\text{SMI},\gamma}(\psi,u,\eta,\widetilde{\eta}\mid\mathcal{D})dud\widetilde{\eta}d\eta$